Raymond is a basketball player who takes four independent free throws with 70% probability of getting a basket on each shot. Let X be the number of baskets Raymond gets. Find the probability that he gets exactly two baskets to 3 decimal places
1
Expert's answer
2020-06-02T19:11:14-0400

We use Bernoulli's formula:

"P(x) = \frac{4!} {2!×(4 - 2)!} ×0.7^2×(1 - 0.7)^{4 - 2} = 6×0.7^2×0.3^2 = 0.265."